Please help me solve this word problem.
Sam's father is 3 times as old as he is. In 6 years, the sum of their ages will be 68. How old is Sam?

Answer by mananth(16949)

Sam's father is 3 times as old as he is. In 6 years, the sum of their ages will be 68. How old is Sam?


We will assume sam's age to be x
Sam's father is three times sam's age = 3x
After six years sam's age will be x+6
and the father's age will be 3x+6
the sum of their ages = 68
x+6+3x+6= 68
4x+12= 68
4x=68-12
4x= 56
x= 56/4
x=14
Sam is 14 years